You are biking at a speed of 18 miles per hour. You are three miles behind your friend, who is biking at a speed of 12 miles per hour. Write and solve an equation to find the amount of time it takes for you to catch up to your friend.
Answer by richwmiller(17219)


Algebraic solution
t*(18-12)=3
t*6=3
t=3/6
t=0.5 hours
check
0.5(18-12)=3
0.5(6)=3
3.0=3
ok
3 12 18 0 0
Gaining speed 18-12= 6 mph
Headstart 3 =3 miles
Time to catch up =0.5 hours